OpenStudy (anonymous):

~Please show work~ What % of 56 is 42?

jhonyy9 (jhonyy9):

56 --- 100 42 --- x ---------- x= (42*100)/56 = 4200/56 =75 so this is 75 %

OpenStudy (nubeer):

(x/100)x56=42 x=75%

OpenStudy (anonymous):

~Please show work~ What % of 48 is 12 ?